# Symmetric quantum states: a review of recent progress

### v2

@@ III.1.2 Symmetric states of qudits
In [22], the above analysis is extended beyond the bipartite case showing that it is still possible to establish a correspondence between a multipartite DS state, $\rho_{DS}^{N}\in\mathcal{S}_{N}^{(d)}$, and a matrix $\tilde{M}_{d}(\rho_{DS})$ arising from partial transposition, although different partitions must be considered. Despite the structure of such matrix is, in general, quite involved, several results have been rigorously proven, which we summarise here: (i) for $N=3$ and $d\leq 4$, PPT is necessary and sufficient for separability; (ii) there is strong analytical and numerical evidence that positivity w.r.t. the largest partition implies also positivity w.r.t. to any other partition, analogously to the case of $N$-qubit DS states; (iii) assuming that condition (ii) is true, then for any $N$ and any local dimension $d$, $\rho_{DS}^{N}$ is separable if and only if $\tilde{M}_{d}(\rho_{DS}^{N})$ is a completely positive matrix.
We conclude this section with a brief survey on entanglement in symmetric states of qudits, beyond the subset of DS states. Starting from the bipartite case, the first examples of PPT-entangled symmetric states were reported in [5], where two methods to generate such states were provided: one relies on embedding entangled states within a Hilbert space of larger dimension, ensuring that both symmetry constraints and PPT condition are satisfied; alternatively, a half-analytical half-numerical method was proposed where bipartite PPT-entangled symmetric states are constructed starting from an $N$-qubit symmetric state which is PPT w.r.t. the partition $\frac{N}{2}:\frac{N}{2}$, while NPT w.r.t. the one of the others (here, $N$ is an even number). With the aid of this technique, the existence of a two-qutrit symmetric bound entangled state was reported. Remarkably, a similar approach, where a multiparty symmetric $N$-qudit DS state is embedded into a bipartite symmetric state of higher local dimension, was adopted also in [22] to relate the entanglement properties of multiparty systems to those of bipartite ones (see Figure 3). This approach also has the advantage of reducing the computational cost needed to assess the separability properties of DS states.
@@
with $\sum_{i}(\lambda_{i}/\lambda^{\prime}_{i})>1$. Here $\lambda^{\prime}_{i}$ is defined as the maximum value such that $\rho_{DS}+\lambda^{\prime}_{i}|D_{ii}\rangle\!\langle D_{jk}|$ is a valid PPT state (i.e., both the state and its partial transposition are positive semidefinite). Notice that when the ranks of the state and its partial transposition are such that $r(\rho,\rho^{\Gamma})=(5,7)$, the corresponding state is edge. Since edge states are extremal within the set of PPT states, understanding their structure is sufficient to characterize the entire PPT set. We conclude this section by noting that, making use of the PPT symmetric extension method [25], the analysis carried out in [22] provides a useful and constructive technique to generate multipartite PPT-entangled symmetric states for arbitrary number of parties, as shown in [23].
@@ III.1.4 Quantification of multipartite entanglement
While for bipartite pure states all entanglement measures coincide, allowing to define with no ambiguity maximally entangled states, moving to multipartite states this is no longer the case and states that maximize one entanglement measure do not necessarily maximize another. Among the various entanglement measures, the geometric measure (GM) of entanglement [33, 34] stands out for its simple definition and operational significance, being related to state discrimination under LOCC [35]. For pure states, the GM quantifies entanglement by measuring the minimal distance between a given state and the set of fully separable states. Mathematically, this requires an optimization procedure over the latter set and, for this reason, even for pure states, calculating the geometric measure is generally a hard task [36]. Quite intuitively, extending the geometric measure to mixed states is even more challenging, since it requires a convex roof construction. This is computationally intractable in general, making exact evaluations unfeasible except in special cases. One such case arises in the study of symmetric states, where remarkable simplifications occur.
@@
The geometric measure (GM) has been employed as a tool to identify candidates for maximally entangled symmetric states in several studies. In particular, [39] analytically derived an upper bound for the GM of pure $N$-qubit symmetric states. Additionally, states exhibiting high geometric entanglement were identified based on an intriguing connection with the Majorana representation, a geometric tool that maps symmetric states to configurations of points on the Bloch sphere. This correspondence was investigated also in [40], where pure symmetric $N$-qubit states with high GM were numerically identified for systems up to $N=20$. The relationship between the Majorana representation and the GM for symmetric states was further explored in [41], where it was used to demonstrate the equivalence between various entanglement measures. More recently, an alternative approach, based on a linear mapping between symmetric subspaces, has been proposed to estimate the amount of entanglement in pure symmetric $N$-qubit states [42] and provide some lower bounds on the GM of a multipartite state.
@@ IV.4 Entanglement certification
where $(\Delta J_{\bm{n}})^{2}\equiv\langle J_{\bm{n}}^{2}\rangle-\langle J_{\bm{n}}\rangle^{2}$ and $J_{\bm{n}}\equiv\bm{n}\cdot\bm{J}$, for a given direction $\bm{n}=(\bm{n}_{1},\bm{n}_{2},\bm{n}_{3})$. Remarkably, in [134], a criterion to detect entanglement in spin-squeezed states was found. Expanding on this result, the relation between spin-squeezing inequalities and entanglement in many-body systems was explored in many other works (see [135] and references therein). In [136, 137, 138], some inequalities were constructed for symmetric $N$-qubit states, providing necessary and sufficient criteria to certify two- and three-qubit genuine multipartite entanglement, while being only necessary for a higher number of qubits.
@@ VII Concluding remarks
Yet, despite this progress, several important open questions remain, pointing to promising avenues for future research. One such challenge is the general characterization of PPT-entangled symmetric states, both in bipartite and multipartite systems. In the case of bipartite diagonal symmetric (DS) states, it has recently been shown that separability is equivalent to a cyclic variation of the complete positivity problem [411], and PPT-entangled DS states have been fully characterized for dimension $d\leq 5$. For multipartite diagonal symmetric states much less is known: while for $N=3$ and $d\leq 4$ a complete solution has been provided in [22], already for $N\geq 4$, the complexity of this task becomes significantly harder, indicating the need for new mathematical tools to address this problem. Studying the properties of non-decomposable entanglement witnesses for symmetric states [412, 27] is also of interest. Aside from their connection with the cone of copositive matrices, advances in this area will led to the discovery of novel properties of positive, but not completely positive, maps. Symmetric states may also play a key role in addressing the conjecture, originally formulated in [413], that no PPT state in $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ can have maximal Schmidt number (that is, equal to $d$). While the conjecture has been proven for $d=3$ [414], extending this result to higher dimension remains a challenging task. A closely related topic is the use of symmetric states to prove the PPT2 conjecture [415, 416]– stating that the composition of two PPT channels is entanglement breaking. For bipartite systems of qudits, this conjecture holds trivially for $d=2$, as a consequence of the PPT criterion. Moreover, in [417, 416], it has been proven for $d=3$, but remains open for $d>3$. When restricting to symmetric states, it was shown in [418] that the conjecture holds true for Choi-type maps, which include bipartite DS states as a special case. However, an analogous result for symmetric, but not DS, states has yet to be established.
@@ Acknowledgements
Carlo Marconi acknowledges support from the European Union - NextGeneration EU, "Integrated infrastructure initiative in Photonic and Quantum Sciences" - I-PHOQS [IR0000016, ID D2B8D520, CUP B53C22001750006]
Guillem Müller-Rigat acknowledges support from: European Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB, I00, project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR" (PRTR-C17.I1), FPI); QUANTERA DYNAMITE PCI2022-132919, QuantERA II Programme co-funded by European Union’s Horizon 2020 program under Grant Agreement No 101017733; Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program; Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, European Commission, European Climate, Infrastructure and Environment Executive Agency (CINEA), or any other granting authority. Neither the European Union nor any granting authority can be held responsible for them (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794, QU-ATTO, 101168628), EU Horizon Europe Program (This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101080086 NeQSTGrant Agreement 101080086 — NeQST); ICFO Internal “QuantumGaudi” project; European Union - NextGeneration EU (PRTR-C17,l1), PPCC-MCIN-L5.
@@
Jordi Tura acknowledge the support received from the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme as well as support received from the European Union’s Horizon Europe research and innovation programme through the ERC StG FINE-TEA-SQUAD (Grant No. 101040729). The views and opinions expressed here are solely those of the authors and do not necessarily reflect those of the funding institutions. Neither of the funding institution can be held responsible for them.
Anna Sanpera acknowledges financial support from Ministerio de Ciencia e Innovación of the Spanish Goverment with funding from European Union NextGenerationEU (PRTR-C17.I1) and by Generalitat de Catalunya and from the European Commission QuantERA grant ExTRaQT (Spanish MICIN project PCI2022-132965), by the Spanish MICIN (project PID2022-141283NB-I00) with the support of FEDER funds, and by the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGeneration EU within the framework of the Digital Spain 2026 Agenda.

### v3

@@ III.1.2 Symmetric states of qudits
The above analysis can be extended beyond the bipartite case showing that it is still possible to establish a correspondence between a multipartite DS state, $\rho_{DS}^{N}\in\mathcal{S}_{N}^{(d)}$, and a matrix $\tilde{M}_{d}(\rho_{DS})$ arising from the partial transposition w.r.t. the largest partition, although, in principle, all possible partitions must now be considered. The structure of such matrix is, in general, quite involved. Nevertheless, there is strong analytical and numerical evidence that positivity w.r.t. the largest partition implies also positivity w.r.t. to any other partition, analogously to the case of $N$-qubit DS states [22]. However, in contrast to the qubit case, already for three qutrit DS states there exist PPTES [23].
We conclude this section with a brief survey on entanglement in symmetric states of qudits, beyond the subset of DS states. Starting from the bipartite case, the first examples of PPT-entangled symmetric states were reported in [5], where two methods to generate such states were provided: one relies on embedding entangled states within a Hilbert space of larger dimension, ensuring that both symmetry constraints and PPT condition are satisfied; alternatively, a half-analytical half-numerical method was proposed where bipartite PPT-entangled symmetric states are constructed starting from an $N$-qubit symmetric state which is PPT w.r.t. the partition $\frac{N}{2}:\frac{N}{2}$, while NPT w.r.t. the one of the others (here, $N$ is an even number). With the aid of this technique, the existence of a two-qutrit symmetric bound entangled state was reported. Remarkably, a similar approach, where a multiparty symmetric $N$-qudit DS state is embedded into a bipartite symmetric state of higher local dimension can be adopted [22] to relate the entanglement properties of multiparty systems to those of bipartite ones (see Figure 3). This approach also has the advantage of reducing the computational cost needed to assess the separability properties of DS states.
@@
with $\sum_{i}(\lambda_{i}/\lambda^{\prime}_{i})>1$. Here $\lambda^{\prime}_{i}$ is defined as the maximum value such that $\rho_{DS}+\lambda^{\prime}_{i}|D_{ii}\rangle\!\langle D_{jk}|$ is a valid PPT state (i.e., both the state and its partial transposition are positive semidefinite). Notice that when the ranks of the state and its partial transposition are such that $r(\rho,\rho^{\Gamma})=(5,7)$, the corresponding state is edge. Since edge states are extremal within the set of PPT states, understanding their structure is sufficient to characterize the entire PPT set. We conclude this section by noting that, making use of the PPT symmetric extension [26], it is possible to generate multipartite PPT-entangled symmetric states for arbitrary number of parties [24].
@@ III.1.4 Quantification of multipartite entanglement
While for bipartite pure states all entanglement measures coincide, allowing to define with no ambiguity maximally entangled states, moving to multipartite states this is no longer the case and states that maximize one entanglement measure do not necessarily maximize another. Among the various entanglement measures, the geometric measure (GM) of entanglement [34, 35] stands out for its simple definition and operational significance, being related to state discrimination under LOCC [36]. For pure states, the GM quantifies entanglement by measuring the minimal distance between a given state and the set of fully separable states (Sep).
$E_{G}(\left|\psi\right\rangle)=1-\max_{\left|\phi\right\rangle\in\text{Sep}}|\langle\phi|\psi\rangle|^{2}.$ (20)
The above expression requires an optimization procedure over the separable set and, for this reason, even for pure states, calculating the geometric measure is generally a hard task [37]. Quite intuitively, extending the geometric measure to mixed states is even more challenging, since it requires a convex roof construction.
$E_{G}(\rho)=\min_{p_{k},\left|\psi_{k}\right\rangle}\sum p_{k}E_{G}(\left|\psi_{k}\right\rangle),$ (21)
where the minimization is taken over all convex decompositions of $\rho$. This is computationally intractable in general, making exact evaluations unfeasible except in special cases. One such case arises in the study of symmetric states, where remarkable simplifications occur.
@@
The GM has been employed as a tool for many things, specially to study multipartite entanglement and to identify candidates for maximally entangled symmetric states and to study their robustness [40]. In particular, [41] analytically derived an upper bound for the GM of pure $N$-qubit symmetric states. Additionally, states exhibiting high geometric entanglement were identified based on an intriguing connection with the Majorana representation, a geometric tool that maps symmetric states to configurations of points on the Bloch sphere. This correspondence was investigated also in [42], where pure symmetric $N$-qubit states with high GM were numerically identified for systems up to $N=20$. The relationship between the Majorana representation and the GM for symmetric states was further explored in [43], where it was used to demonstrate the equivalence between various entanglement measures. More recently, an alternative approach, based on a linear mapping between symmetric subspaces, has been proposed to estimate the amount of entanglement in pure symmetric $N$-qubit states [44] and provide some lower bounds on the GM of a multipartite state. For 3 qubits highly symmetric states, such as the GHZ-symmetric and W states (and mixtures of those), the 3-tangle measure[45] has also been computed [46, 47].
@@ IV.4 Entanglement certification
where $(\Delta J_{\bm{n}})^{2}\equiv\langle J_{\bm{n}}^{2}\rangle-\langle J_{\bm{n}}\rangle^{2}$ and $J_{\bm{n}}\equiv\bm{n}\cdot\bm{J}$, for a given direction $\bm{n}=(\bm{n}_{1},\bm{n}_{2},\bm{n}_{3})$. Remarkably, in [139], a criterion to detect entanglement in spin-squeezed states was found. Expanding on this result, the relation between spin-squeezing inequalities and entanglement in many-body systems was explored in many other works (see [140] and references therein). In [141, 142, 143], some inequalities were constructed for symmetric $N$-qubit states, providing necessary and sufficient criteria to certify two- and three-qubit genuine multipartite entanglement, while being only necessary for a higher number of qubits. Nevertheless, other approaches to detect GME around qubit Dicke states have also been developed [144], showing that in the vicinity of Dicke states there always exists GME.
@@ VII Concluding remarks
Yet, despite this progress, several important open questions remain, pointing to promising avenues for future research. One such challenge is the general characterization of PPT-entangled symmetric states, both in bipartite and multipartite systems. In the case of bipartite diagonal symmetric (DS) states, it has recently been shown that separability is equivalent to a cyclic variation of the complete positivity problem [417], and PPT-entangled DS states have been fully characterized for dimension $d\leq 5$. For multipartite diagonal symmetric states much less is known. While for N=3 and d=3 there exist PPT-Bound DS states [23], the complexity of this task becomes significantly harder as the number of parties or local Hilbert dimension increases [22], indicating the need for new mathematical tools to address this problem. Studying the properties of non-decomposable entanglement witnesses for symmetric states [418, 28] is also of interest. Aside from their connection with the cone of copositive matrices, advances in this area will led to the discovery of novel properties of positive, but not completely positive, maps. Symmetric states may also play a key role in addressing the conjecture, originally formulated in [419], that no PPT state in $\mathbb{C}^{d}\otimes\mathbb{C}^{d}$ can have maximal Schmidt number (that is, equal to $d$). While the conjecture has been proven for $d=3$ [420], extending this result to higher dimension remains a challenging task. A closely related topic is the use of symmetric states to prove the PPT2 conjecture [421, 422]– stating that the composition of two PPT channels is entanglement breaking. For bipartite systems of qudits, this conjecture holds trivially for $d=2$, as a consequence of the PPT criterion. Moreover, in [423, 422], it has been proven for $d=3$, but remains open for $d>3$. When restricting to symmetric states, it was shown in [424] that the conjecture holds true for Choi-type maps, which include bipartite DS states as a special case. However, an analogous result for symmetric, but not DS, states has yet to be established.
@@ Acknowledgements
Carlo Marconi acknowledges support from the European Union - NextGeneration EU, "Integrated infrastructure initiative in Photonic and Quantum Sciences" - I-PHOQS [IR0000016, ID D2B8D520, CUP B53PGC22001750006]
Guillem Müller-Rigat acknowledges support from: European Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB, I00, project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR" (PRTR-C17.I1), FPI); QUANTERA DYNAMITE PCI2022-132919, QuantERA II Programme co-funded by European Union’s Horizon 2020 program under Grant Agreement No 101017733; Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program; Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794, QU-ATTO, 101168628), EU Horizon Europe Program (This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101080086 NeQSTGrant Agreement 101080086 — NeQST); ICFO Internal “QuantumGaudi” project; European Union - NextGeneration EU (PRTR-C17,l1), PPCC-MCIN-L5.
@@
Jordi Tura acknowledge the support received from the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL programme as well as support received from the European Union’s Horizon Europe research and innovation programme through the ERC StG FINE-TEA-SQUAD (Grant No. 101040729).
Anna Sanpera acknowledges financial support from Ministerio de Ciencia e Innovación of the Spanish Goverment with funding from European Union NextGenerationEU (PRTR-C17.I1) and by Generalitat de Catalunya and from the European Commission QuantERA grant ExTRaQT (Spanish MICIN project PCI2022-132965), by the Spanish MICIN (project PID2022-141283NB-I00) with the support of FEDER funds, and by the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGeneration EU within the framework of the Digital Spain 2026 Agenda. The views and opinions expressed here are solely those of the authors and do not necessarily reflect those of the funding institutions. Neither of the funding institution can be held responsible for them.